Four microprocessors are randomly selected from a lot of 100 microprocessors among which 10 are defective. Find the probability of obtaining no defective microprocessors.


1
Expert's answer
2021-11-10T09:21:00-0500

The probability of obtaining no defective microproccessors is equal to probability that every obtained microprocessor is good

Let A be "all microprocessors are good", "A{\scriptscriptstyle n}" be "n-th microprocessors is good", then

"P(A{\scriptscriptstyle 1})= {\frac {90} {100}}"

"P(A{\scriptscriptstyle 2})= {\frac {89} {99}}"

"P(A{\scriptscriptstyle 3})= {\frac {88} {98}}"

"P(A{\scriptscriptstyle 4})= {\frac {87} {97}}"

Since all probabilities were calculated considering the changes in the amount of the detais after each pick, then

"P(A) =P(A{\scriptscriptstyle 1})*P(A{\scriptscriptstyle 2})*P(A{\scriptscriptstyle 3})*P(A{\scriptscriptstyle 4})={\frac {90*89*88*87} {100*99*98*97}}=0.6516" (approximately)